OpenStudy (anonymous):

Television sets. What does it mean to refer to a 20 in TV set or 25 in TV set? Such units refer to the diagonal of the screen. A 30 in TV set also has a width of 24 inches. What is its height? What is the height of a 30 in TV? ___ inches

OpenStudy (anonymous):

TVs are measured by the diagonal of their screens

OpenStudy (anonymous):

phytagorean theorem says diagonal D,W=width,H=height D^2=W^2+H^2 30^2 =24^2 +H^2 30^2 -24^2 =H^2 324=H^2 sqrt324=H 18=H height ans